OpenStudy (kaylala):

Find the value of the following using the concepts of both addition and subtraction formulas and reference angle: cos 195

OpenStudy (anonymous):

cos (195) = cos (180 + 15) = - cos (15) Do you know how to find cos (15)?

OpenStudy (kaylala):

cos (45-30) = cos 15? @MonoPolyRoly

OpenStudy (anonymous):

Yes.

OpenStudy (kaylala):

what to do next @MonoPolyRoly

OpenStudy (anonymous):

Use the compound angle formula cos(A-B) = cosAcosB + sinAsinB.

OpenStudy (kaylala):

= ((sqrt6)/4) + ((sqrt2)/4) ? @MonoPolyRoly

OpenStudy (anonymous):

That's cos(15). Note that cos(195) = -cos(15). So...

OpenStudy (kaylala):

what happens? why the negative? @MonoPolyRoly

OpenStudy (anonymous):

I did explain to you why cos(195) = -cos(15) in my first comment here. Please scroll up and check it. :)

OpenStudy (kaylala):

i mean where did you get the - from what derivation? @MonoPolyRoly

OpenStudy (kaylala):

= - ((sqrt6)/4) + ((sqrt2)/4) is that it? @MonoPolyRoly

OpenStudy (anonymous):

It is. Also, cos (195) = cos (180 + 15) = - cos (15). Not that cos (180 + x) = -cos(x)